OpenStudy (anonymous):

The expression 2/(x-2)(x+3) is undefined when x is equal to: a.0, b:2 or -3 c. -2 or 3 or d. -1

OpenStudy (anonymous):

b:2 or -3 when x=2, the expression 2/(x-2)(x+3) becomes 2/(2-2)(x+3)=2/(0)(x+3)=undefined similarly when x=-3,the expression 2/(x-2)(x+3) becomes 2/(x-2)(-3+3)=2/(x-2)(0)=undefined